OpenStudy (anonymous):

Expand the expression. logc 5/8

OpenStudy (anonymous):

log5-log8

OpenStudy (anonymous):

according to the rule logb(m/n) = logb(m) – logb(n) logc(5/8) = logc(5) – logc(8)

OpenStudy (anonymous):

@marunadl can you tell me all the rules to expand and condense a logarithmic expression?

OpenStudy (anonymous):

http://www.purplemath.com/modules/logrules.htm